OpenStudy (anonymous):

y = 2x - 2 12x - y = 22 A. (2, 2) B. (2.4, 2.8) C. (- 2, - 6) D. (- 2.4, - 6.8) y = 2x - 2 12x - y = 22 A. (2, 2) B. (2.4, 2.8) C. (- 2, - 6) D. (- 2.4, - 6.8)

OpenStudy (texaschic101):

y = 2x - 2 sub 2x - 2 in for y in your other equation 12x - y = 22 12x - (2x - 2) = 22 12x - 2x + 2 = 22 12x - 2x = 22 - 2 10x = 20 x = 2 y = 2x - 2 y = 2(2) - 2 y = 4 - 2 y = 2 x = 2 and y = 2

OpenStudy (anonymous):

c